What is f(x)=3x+2 and g(x)=2x+5. Find fog(5)
Answer by Theo(13342)

f(x) = 3x+2

g(x) = 2x+5

fog(x) = f(g(x))

g(5) = 2*5+5 = 15

fog(5) = f(g(5)) = f(15)

f(15) = 3*15+2 = 47

your answer is that fog(5) = 47